OpenStudy (faman39):

A container states that it should be stored at a temperature above 15 c . Convert the temperature to degrees Fahrenheit. Round to the nearest whole number, if necessary. 9 F 31 F 40 F 59 F

OpenStudy (faman39):

@ganeshie8 can you able to help me out bit?

OpenStudy (faman39):

@mark_o. can you able to help me bit?

OpenStudy (anonymous):

degrees F=1.8 C +32 or F= 9/5 C +32 try to convert them first and lets see

OpenStudy (faman39):

ah, actually i forget the process, i m so sorry, i hate that habits of mine :(

OpenStudy (anonymous):

thats ok, it would be better if you keep your notes, then review them once you needed them

OpenStudy (faman39):

so ok, i think it would be 9 but we do something subtract or something

OpenStudy (faman39):

that we need to do, thats all i remember

OpenStudy (anonymous):

ok the problem says the container need to be above 15C=59F, so it is supposed to cool up if we are talking about container to cool up not to heat up,

OpenStudy (anonymous):

so its supposed to be 59F or less........

OpenStudy (faman39):

less, so 9?

OpenStudy (faman39):

no 40

OpenStudy (anonymous):

the freezing point of water is 0C= 1.8(0)+32= 32F

OpenStudy (anonymous):

ill say 31F is closer to 32F

OpenStudy (faman39):

ok thank you sooo much! :) i will try to review as u said i will do it good

OpenStudy (faman39):

Sorry for giving so much trouble for it, actually i m really dumb abt math

OpenStudy (anonymous):

no thats ok, it just takes practice,,you just need to practice and solved more problems,,,

OpenStudy (faman39):

Thank you so much @mark_o. :) ok i m trying it to convert

OpenStudy (anonymous):

if F=9/5 C +32 then (F-32) *5/9 =C how much degree C is 31F? try to changed that

OpenStudy (faman39):

ok are we need to divide 5/6 first, right?

OpenStudy (anonymous):

ok YW, if you have more prob there try to post them, and i will try to help you out as much as i can,,, enjoy solving now...good luck :D

OpenStudy (anonymous):

heres the formula, just plug them in (F-32) *5/9 =C

OpenStudy (faman39):

ok thanks alot helping for so much ! :)

OpenStudy (anonymous):

(F-32) *5/9 =C (31-32) *5/9 =C C=-5/9= - 0.556 approx -1 C so 31F=-1C

OpenStudy (faman39):

ok thank you so much @mark! :) sorry for bothering, i m really dumbest person in math

OpenStudy (anonymous):

ok enjoy solving ...:D

OpenStudy (anonymous):

no prob....post them anytime and send me note and i will read them ..lol ...good luck now

OpenStudy (faman39):

:) thanks alot